sps ls1 planning meeting n014 05 06 13 n.
Download
Skip this Video
Loading SlideShow in 5 Seconds..
SPS LS1 Planning meeting ( N014) 05 /06/13 PowerPoint Presentation
Download Presentation
SPS LS1 Planning meeting ( N014) 05 /06/13

Loading in 2 Seconds...

play fullscreen
1 / 20
charles-crawford

SPS LS1 Planning meeting ( N014) 05 /06/13 - PowerPoint PPT Presentation

86 Views
Download Presentation
SPS LS1 Planning meeting ( N014) 05 /06/13
An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author. While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.

- - - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript

  1. SPS LS1Planning meeting (N014)05/06/13 David Mcfarlane SPS.Technical-Coordination@Cern.ch

  2. Topics • Lifts and Monte-charges. • Magnet change campaign • Works affecting access. • TT10 Update. • Valve replacement on the SPS ring-main • Planning. • False floors open in the BAs • E.C.Rs. • Access rights and safety equipment • Important Information. • A.O.B SPS.Technical-Coordination@Cern.ch

  3. Lift Maintenance June/July SPS.Technical-Coordination@Cern.ch

  4. Works on the Monte-Charges Hi David, For the meeting this afternoon , I would like you ask to the users if AS Ascenseurs could arrest the monte-charges to proceed at safety works.It needs 2 days per monte-charge. These works have been requested by J.Pedersen following a important safety defiency during the usual maintenance. AS Ascenseurs would prefer the dates below but all is questionable: Week 26 fromJune 24 to 28 Week 27 from July 03 to 05 Week 28 from July 08 to 12 Best regards. Nicolas SPS.Technical-Coordination@Cern.ch

  5. Magnet change campaign(week 24) SPS.Technical-Coordination@Cern.ch

  6. Magnet change campaign(week 25) SPS.Technical-Coordination@Cern.ch

  7. Works affecting access SPS.Technical-Coordination@Cern.ch

  8. TT10 update • Presentation by Richard Morton SPS.Technical-Coordination@Cern.ch

  9. Valve Replacement ‘Boucle SPS’ – LS1 • Presentation by Bill Bannister SPS.Technical-Coordination@Cern.ch

  10. SPS LS1 schedule Open Access : 27/03/13 Close Access : 27/06/14 Beam off : 16/02/13 Beam on: 13/10/14 Beam on Cool down/RP survey and Magnet tests Cool down NO ACCESS Kicker Survey General Maintenance Survey General Maintenance Survey Survey Cables EPC works RF upgrade works RF upgrade works EPC works Survey Cables Survey Kicker Survey 18 KV Civil Engineering in TT10 Fibres Magnets Magnets Magnets Magnets Magnets Magnets Kicker Cables Magnets Magnets 18 KV General maintenance General Maintenance Survey General maintenance CrabCavities General Maintenance Cooling tower 863 Vacuum Septa works Septa works Vacuum Vacuum Ring-main Ring-main Ring-main Ring-main Ring-main Fibres Ring-main Ring-main General Maintenance 18 KV General Maintenance Vacuum Prep work in LSS1 Vacuum Survey Horizontal Vertical smoothing RF Upgrade works Cables Cool down Limited ACCESS UA9 works TT20 pipework 18 KV Fibres CV maintenance CV maintenance Civil Engineering CV maintenance 18 KV Fibres Irradiated Cables Cables Cables Cables Fibres Cables 18 KV EL Studies Civil Engineering TT10 Consolidation CV studies Cables Septa replacement Civil Engineering Civil Engineering Civil Engineering CrabCavities RF Upgrade works Civil Engineering RF Upgrade works Re-install equipment removed for cable works RAL CV works Magnets Fibres Cables Magnets Magnets Magnets Magnets Magnets Magnets Magnets EPC Commissioning tests Vacuum DSO tests/HV tests Cold check out & Start-up Protons to the North Area Protons to LHC Ions to NA61 CV Controls upgrade Re-route Cabling

  11. False floors lifted in Surface BAs • Below is a table showing the dates that EL will be pulling cables at each point of the SPS. • During these periods large sections of the false floor in these surface buildings will be lifted. • Any works in the BA’s that involve lifting of the false floors MUST be reported to the Coordination team • If you require access to racks or other equipment in the surface buildings during these periods you must check with planning first. SPS.Technical-Coordination@Cern.ch

  12. Engineering Change Requests • ECR link • http://sps.web.cern.ch/SPS/quality.htm • ECR Procedure • http://sps.web.cern.ch/SPS/files/quality/ECR_note_gestion_rev02.pdf • ECR Template • http://sps.web.cern.ch/SPS/files/quality/ECR_SPS_formulaire_FR.pdf • All ECRs must be submitted to Jerome Axensalva (BE/OP) AND myself. SPS.Technical-Coordination@Cern.ch

  13. Current ECR’s for LS1 https://edms.cern.ch/nav/P:CERN-0000076703:V0/P:CERN-0000096067:V0/TAB3 • Removal of AEWB Pick-up from position 31932 to 31502 • (EDMS No. 1245361) RELEASED • Removal of a BPWA Pick-up from position 31931 to 31101 • (EDMS No. 1245393)RELEASED • Installation of a new Wideband Kickers system in position 32173 • (EDMS No. 1245400)RELEASED • Installation of a new HBW Damper Pick-up in place of BPWA 31931 & AEWB 31932 • (EDMS No. 1245401)RELEASED • Replacement of 4 old BPA and 2 BPD by 6 BPCE in LSS1 & LSS2 • (EDMS No. 1246559)RELEASED • Reinstallation of pump and gauges (Pirani + Penning) in 11993 (vacuum sector 150) • (EDMS No. 1249305)RELEASED • Modification LSS1 sectorization (new VVSF + displacement VVSB 11772) • (EDMS No. 1249307 v.2)RELEASED • Installation of 3 BCT in LSS5 for North Area protection during Ions operations in SPS. • (EDMS No. 1075945) UNDER IMPLEMENTATION SPS.Technical-Coordination@Cern.ch

  14. Current ECR’s for LS1 https://edms.cern.ch/nav/P:CERN-0000076703:V0/P:CERN-0000096067:V0/TAB3 • Transfer of the Fast Beam Current Transformers (BCTFR.31451) for the SPS circulating beam from LSS3 to LSS5. • (EDMS No. 1221341) RELEASED • SPS Beam Position Monitoring Prototyping Setup in LSS4. • (EDMS No. 1099630)RELEASED • Installation de 4 moniteurs de position (2 BPH, 2 BPV), dédiés au système Damper. • (EDMS No. 1251693)RELEASED • Installation d'un nouveau moniteur de position de type BPCR en LSS2 au 22105. • (EDMS No. 1251695) RELEASED • Displacement of Door PPG2145 separating the LHC and the SPS in the TI2 tunnel. • (EDMS No. 1253832)RELEASED • Interlocking of Injection Tunnel upstream areas with LHC powering. • (EDMS No. 1257949)RELEASED • Move of BPH.31998 to BPH.42173. • (EDMS No. 1272718) RELEASED • Optical Fibre Racks. • (EDMS No. 1273766) RELEASED SPS.Technical-Coordination@Cern.ch

  15. Current ECR’s for LS1 https://edms.cern.ch/nav/P:CERN-0000076703:V0/P:CERN-0000096067:V0/TAB3 • Removal of old fixed TMP vacuum groups in 11655-11691. • (EDMS No.1273232)RELEASED • BYBPM racks under the SPS dipole magnets as part of the Upgrade of Beam Position Monitors electronics (MOPOS). • (EDMS No. 1282651)UNDER IMPLEMENTATION • Installation de Coffrets d'Alimentation en LSS1 et TT10. • (EDMS No. 1282860)RELEASED • Installation d'un rack de jonction RX11695=TS1 au BA1, afin de limiter les longueurs des câbles à changer lors d'un futur remplacement. • (EDMS No. 1288607)UNDER APPROVAL SPS.Technical-Coordination@Cern.ch

  16. Access rights and safety equipment • Everyone who enters the SPS MUST have the following: • Completed the correct courses. • Safety course levels 1,2 and 3. • HabilitationElectrique. • Radiological Protection. • Self Rescue Mask Training. • The correct access authorisation (via EDH). • A valid approved IMPACT. (with their name included on this IMPACT) • They must know the number of this IMPACT. • Safety shoes. • Biocell. • Electronic dosimetre. • Passive dosimetre. • Helmet plus Light. • Please note that the medical service no longer contact you when your medical is about to expire!!!!!!! • You will receive an e-mail 2 months before informing you that some of your access rights will soon expire. • It is up to you to contact the medical service to arrange a medical. SPS.Technical-Coordination@Cern.ch

  17. Important information • Useful Contact Numbers • SPS Coordination Team • David MCFARLANE- 164247 • Franck BAIS – 165467 • James RIDEWOOD - 160590 • Control Room • SPS CCC – 77500 • TI CCC – 72201 • CSA - 62511 • RP • Christophe Tromel – 163199 • Angelito Herve – 163168 • Nadine Conan (BA1) – 160641 • Kurt Weiss (BA1) -160759 • Coordination Safety • Emmanuel Paulat – 163870 • John Etheridge - 164647 SPS.Technical-Coordination@Cern.ch

  18. A.O.B. • One of the old electric chariots was found broken down and abandoned in the tunnel at BA6. If you have a problem with one of the chariots you must contact us. DO NOT just leave it where it is and not tell anyone. • There are currently 55 IMPACTs that are “LATE”. If you have completed your work then you need to close the related IMPACT. SPS.Technical-Coordination@Cern.ch

  19. ECA4 - Access to platforms Access to this area seems to be blocked - Is it for radiation or general safety issues? - What are access conditions with/without beam? - Who is responsible for access conditions here? SPS.Technical-Coordination@Cern.ch

  20. SPS coordination website http://sps.web.cern.ch/SPS/ http://cern.ch/LS1planning \\cernhomeJ.cern.ch\J\jocksoft\Public\LS1 Next meeting 19/06/2013 892-1-D20 SPS.Technical-Coordination@Cern.ch